OpenStudy (anonymous):

Express the exact value of csc 60 degrees, with a rational denominator.

OpenStudy (anonymous):

cscx=1/sinx

OpenStudy (anonymous):

= 1 / (sqrt3 / 2) = 2 / sqrt 3

OpenStudy (anonymous):

60 degrees = pi/3 in unit circle

OpenStudy (anonymous):

2 / sqrt 3=2sqrt3/3 that's your answer